OpenStudy (anonymous):

can someone help me with these 2 questions.

OpenStudy (anonymous):

what is the slope-intercept form of the linear equation whose graph passes through (-2,4) and is parallel to the graph 3x+y=6? a) y =-3x - 2 b) y =-3x + 10 c) y = 3x - 2 d) y = 3x + 10 which linear equation represents a line that passes through (3,-7) and has a slope of -4? a) y = -4x-19 b) y = -4x-5 c) y = -4x+5 d) y = -4x+19

OpenStudy (fortytherapper):

For the first one, you have to get y by itself. If you want to get 3x to the other side of the equal sign, what would you do?

OpenStudy (anonymous):

umm subtract 3x from both sides? idk...